Determine the pH for the following solutions: Part A [OH-] = 2.7x10-7M Express your answer using two decimal places. ANSWER:Problem 10.92 Calculate the [H3O+] and [OH-] for a solution with the following pH values: Part A pH = 10.5 Express your answePart E pH = 1.00 Express your answers using two significant figures separated by a comma. ANSWER: [H3O+],[OH-] = M
Answer #1

[H+] [OH-] = 10-14

pH = - log [H+]

Part A

[OH-] = 2.7 x 10-7 M

[H+] = 10-14 / 2.7 x 10-7

[H+] = 3.7 x 10-8

pH = - log [3.7 x 10-8]

pH =7.43

Part B

[H3O+] = [H+] = 4.7 x 10-3 M

pH = - log [4.7 x 10-3 ]

pH = 2.32

Part C

[H3O+] = [H+] = 2 x 10-4 M

pH = - log [ 2 x 10-4 ]

pH = 3.69

Part D

[OH-] = 8.0 x 10-9 M

[H+] = 10-14 / 8.0 x 10-9

[H+] = 1.25 x 10 -8

pH = - log [ 1.25 x 10 -8 ]

pH = 7.9